OpenStudy (anonymous):

Use the quadratic formula to solve x2 + 8x = –2

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

\[x^2 + 8x = –2\] --add +2 to both sides. \[x^2 + 8x + 2 = 0\] now use the quadratic formula.

OpenStudy (anonymous):

x2+8x=-2 +2 +2 x2+8x+2=0 a=1 b=4 c=2

OpenStudy (anonymous):

where: a = 1 b = 8 c = 2 plug them into the formula.

OpenStudy (anonymous):

do the 8 turn positive

OpenStudy (anonymous):

the 8 stays positive.

OpenStudy (anonymous):

but b = 8 not 4... you put b = 4?

OpenStudy (anonymous):

I got 8+/- {8}/2

OpenStudy (anonymous):

wait that's wrong

OpenStudy (anonymous):

im stuck

OpenStudy (anonymous):

what are you stuck on? Any step in particular?

OpenStudy (anonymous):

yes it the step after I got 8+/-(64-8 ____________ 2 8+/-(56 _________ 2

OpenStudy (anonymous):

\[x = \frac{ -(8) \pm \sqrt{(8)^2 - 4(1)(2)} }{ 2(1) }\] \[x = \frac{ -8 \pm \sqrt{64 - 8} }{ 2 }\] \[x = \frac{ -8 \pm \sqrt{56} }{ 2 }\] \[x = \frac{ -8 \pm 2 \sqrt{14} }{ 2 }\] Can you simplify from here? Any questions just ask away :D

OpenStudy (anonymous):

i may have to watch a video on youtube....quadratic's are one of my weak topic's

OpenStudy (anonymous):

@some_someone how'd you get to the 14? and do the -8 turn into a positive 8?

OpenStudy (anonymous):

\[\sqrt{56} = 2\sqrt{14}\]